OpenStudy (anonymous):

Find a formula for tn so that the first four terms are 15, 25, 35, 45. A. tn = 15n B. tn = 10n - 5 C. tn = 10n + 5 D. tn = 15n - 5

OpenStudy (anonymous):

The answer would be C. N is the number of the term you are trying to find, so using the equation in choice C, tn = 10n +5, you can fill in the number of the term you are trying to find in for n. First term, (1) = 10(1)+5 = 15 10(2)+5 = 25 10(3)+5=35 10(4)+5=45 The fifth term would then be 55.

OpenStudy (anonymous):

im so glad you explained it because for some reason i never get it

OpenStudy (anonymous):

Find the first four terms of the given sequence tn = (-1)n(n). A. -1, -2, -3, -4 B. -1, 2, -3, 4 C. 1, ½, 2/3, ¾ D. 1, 2, 3, 4 be my helping buddy im taking a test for summer school